OpenStudy (anonymous):

Solve the system by subsitution 6x+3y=-6 2x+y=-2

OpenStudy (anonymous):

This is a fun one. You sub in the values like so: 6x + 3y = -6 Isolate y 3y = -6 - 6x Divide it all by 3 y = -2 - 2x

OpenStudy (anonymous):

So now you have your y value. Let's plug it into the first equation.

OpenStudy (anonymous):

6x + 3(-2 -2x) = -6

OpenStudy (anonymous):

6x -6 - 6x = -6 (little satanic. derp) Combine like terms: -6 = -6

OpenStudy (anonymous):

You can test it again with the lower equation: 2x+y=-2 2x + (-2 - 2x) = -2 2x - 2 - 2x = -2 -2 = -2

OpenStudy (anonymous):

I don't see that answer on there.. It's either, x=0, y=-2 infinite solutions x=-1, y=0 no solution

OpenStudy (anonymous):

So it's no solution?

OpenStudy (anonymous):

I think it's infinite solutions.

OpenStudy (anonymous):

No unique solution. The first equation is 3 times the second.

OpenStudy (anonymous):

I forget what it means, either none, or infinite.

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (anonymous):

No problem.